### 31 Mar 2026 v1

@@ 10.2.4 Two-Stage Branch
Including the contrastive-only results clarifies the role of the reconstruction stage. Under linear probing, the best two-stage models remain the strongest STRADAViT configurations overall, but the extra reconstruction stage helps mainly on MiraBest and LoTSS DR2; RGZ DR1 is already well served by contrastive-only training. Relative to the strongest off-the-shelf DINOv2 baseline, the best two-stage probe result remains below DINOv2 on MiraBest but exceeds it on LoTSS DR2 and RGZ DR1.
@@ 10.3 Selected STRADAViT Model: Summary and Classwise Analysis
Table 15 summarizes the delta comparisons for both transfer settings. Relative to ViT-MAE, the gains are positive on all three datasets under linear probing and remain positive on MiraBest and RGZ DR1 under fine-tuning. Relative to the strongest DINOv2 baseline, the gains are selective: positive on LoTSS DR2 and RGZ DR1 under probing, and on MiraBest and RGZ DR1 under fine-tuning. The smaller fine-tuning deltas indicate that supervised adaptation closes part of the representation gap observed under linear probing. Taken together with the targeted DINOv2 ablation, these results support the ViT-MAE-based two-stage checkpoint as the primary STRADAViT release: it remains competitive with the stronger DINOv2-based alternative while retaining a substantially lower token count and lower downstream cost. Figure 4 decomposes these aggregate results at class level.
@@
Under linear probing, shown in Figure 4(a), the clearest classwise advantage appears on RGZ DR1: the selected model strengthens all six diagonal entries, reaching 92%, 81%, 78%, 85%, 72%, and 87% from 1C1P through 3C3P. The gain is therefore distributed across the full label space rather than concentrated in a single easy row. MiraBest is also favorable in this comparison, with the selected model reaching 68% on FR I and 72% on FR II, ahead of both ViT-Mae and DINOv2 (Registers). LoTSS DR2 remains mixed: the selected model leads on FR I, FR II, and Spiral, but the Hybrid and Relaxed-double rows remain strongest for ViT-Mae.
Under full fine-tuning, shown in Figure 4(b), the classwise margins are smaller. On MiraBest, all three models are tied at 70% on FR I, while the selected model matches the strongest FR II row at 75%. LoTSS DR2 remains the clearest limitation: the selected model is strongest only on FR II (88%) and trails at least one baseline on the remaining rows. RGZ DR1 remains partly favorable, but the gain is more localized than under probing: the selected model matches the best 1C1P row at 93% and leads on 1C2P and 2C2P (90% and 94%), whereas DINOv2 (Registers) remains stronger on 1C3P, 2C3P, and 3C3P. The fine-tuning deltas are therefore associated with selective classwise changes rather than with a uniform shift of the full matrix.
@@ 10.4 Overall Interpretation and Limitations
The branch comparison indicates that additional stages are not uniformly beneficial across settings. Reconstruction-only continued pretraining is the weakest of the three branches and often underperforms the ViT-MAE starting point. Masked reconstruction alone is therefore not sufficient to improve transfer in this setting. Introducing contrastive learning changes the picture substantially: contrastive-only training already produces large gains over phase 1 alone, especially under linear probing. Across the branch comparison, the two-stage branch is the most consistently top-performing STRADAViT variant, even though a small number of settings—most notably fine-tuned RGZ DR1—are matched or marginally exceeded by contrastive-only training.
@@
The comparison against the actual initialization is informative in two respects. First, relative to ViT-MAE, the strongest two-stage models improve on all linear-probe settings and on fine-tuned MiraBest and RGZ DR1. Second, the branch comparison shows that most of this benefit comes from the contrastive objective, while the reconstruction stage contributes a smaller, dataset-dependent benefit. Phase 1 is best interpreted as a domain-adaptive warm start or regularizer for the subsequent contrastive stage, not as the main source of representation quality by itself. Two factors appear most influential: contrastive losses and ROI-aware view generation. The former better fits a setting in which many negatives are easy background cases while a smaller set are genuinely confusable; the latter makes both contrastive learning and masked reconstruction workable on sparse radio astronomy cutouts. This difference between frozen-feature gains and end-to-end gains is explicit in Table 15 and Figure 4: linear-probe improvements are larger and more consistent across datasets, whereas fine-tuning gains are smaller, more localized to specific classes, and change sign on LoTSS DR2. The classwise comparisons therefore indicate broader representation-level gains under probing than under end-to-end adaptation.
@@
Across the contrastive objectives, InfoNCE provides a baseline but rarely defines the strongest probe results. Soft-HCL is the most consistent choice on LoTSS DR2 and in many register-based settings, whereas HCL tends to give the best MiraBest results and remains strongest on RGZ DR1 fine-tuning. Register-token effects are also more favorable inside the STRADAViT pipeline than in the off-the-shelf DINO baselines, especially on LoTSS DR2 and RGZ DR1 under probing, although the gain is not uniform across every branch and dataset. The DINOv2-Base(R)-initialized ablation is consistent with this pattern: it indicates portability of the contrastive adaptation recipe to a stronger starting point, but the resulting gains remain small and dataset-dependent, with the same LoTSS DR2 limitation that appears across the adapted models more generally.
@@
The view generator is strongly object-centric. This property reduces empty crops and supports training on compact or moderately extended sources, but it also biases the pipeline toward RGZ-style views. LoTSS DR2 and MiraBest contain larger-scale extended structure that benefits from larger-view sampling designed for extended source morphology in radio astronomy.
@@ 11 Conclusions and Future Work
This paper studies how to adapt self-supervised ViT continued pretraining to sparse, heterogeneous radio astronomy imaging. The main result is that masked reconstruction alone is insufficient, contrastive learning is the main source of transfer gain, and the two-stage reconstruction-to-contrastive recipe is the most consistent STRADAViT configuration overall. The targeted DINOv2 ablation indicates that this adaptation recipe is not specific to the ViT-MAE starting point, but the ViT-MAE-based two-stage checkpoint remains the preferred STRADAViT release because it combines competitive transfer with substantially lower token count and lower downstream cost than the DINOv2-based alternative. In that sense, STRADAViT is best interpreted as a domain-adapted starting point for radio astronomy vision backbones: it provides a better practical replacement for out-of-the-box ViTs in frozen-backbone transfer, and remains favorable in most fine-tuning settings without implying uniform superiority in every dataset or class. We release this ViT-MAE-based checkpoint on Hugging Face [^1] https://huggingface.co/ISSA-ML/stradavit-base. Code, configuration files, and dataset-construction scripts are available from the corresponding author upon reasonable request; where direct redistribution of underlying data products is not possible, the reconstruction pipeline and source references can be provided.
Future work should test stronger teacher–student pretraining schemes, especially DINO-style variants built on anchored radio astronomy-aware view generation, together with larger-view regimes that preserve extended emission more consistently in pretraining. Further iterations of the on-the-fly data-curation and view-generation pipeline are also warranted, especially through augmentation and sampling schemes that better match the heterogeneity of the downstream datasets rather than remaining predominantly RGZ-like, while maintaining the general recipe presented here. Such refinements should improve the cross-dataset balance of the learned representations and are therefore a plausible route to more uniform downstream gains. The strong linear-probe gains also motivate evaluation in fixed-backbone settings, including detection and segmentation pipelines in which STRADAViT can serve as a domain-adapted feature extractor in place of weaker non-specialized ViT backbones. In parallel, the stable downstream gains observed under full fine-tuning justify further evaluation of task-specific models built directly on top of STRADAViT. Additional priorities are broader cross-survey robustness tests, explicit out-of-domain transfer, and repetition of the same controlled pipeline from stronger off-the-shelf initializations.
@@ Acknowledgements
The authors acknowledge Xjenza Malta for funding this study under the Research Excellence Programme of 2024 [REP-2024-19]. The authors thank Alessio Magro and Kristian Grixti, Institute of Space Sciences and Astronomy, University of Malta, for technical support and for maintaining the GPU computing infrastructure used for the training and evaluation runs in this study.

### 08 Apr 2026 v3

@@ 10.2.4 Two-Stage Branch
Including the contrastive-only results clarifies the role of the reconstruction stage. Under linear probing, the best two-stage models remain the strongest STRADAViT configurations overall and are therefore the basis of the selected release checkpoint, but the extra reconstruction stage helps mainly on MiraBest and LoTSS DR2; RGZ DR1 is already well served by contrastive-only training. Relative to the strongest off-the-shelf DINOv2 baseline, the best two-stage probe result remains below DINOv2 on MiraBest but exceeds it on LoTSS DR2 and RGZ DR1.
@@ 10.3 Selected STRADAViT Model: Summary and Classwise Analysis
Table 15 summarizes the delta comparisons for both transfer settings. Relative to ViT-MAE, the gains are positive on all three datasets under linear probing and remain positive on MiraBest and RGZ DR1 under fine-tuning. Relative to the strongest DINOv2 baseline, the gains are selective: positive on LoTSS DR2 and RGZ DR1 under probing, and on MiraBest and RGZ DR1 under fine-tuning. Together with the targeted DINOv2 ablation, these results support selection of the ViT-MAE-based two-stage checkpoint as the primary STRADAViT release: it remains competitive with the stronger DINOv2-based alternative while retaining a substantially lower token count and lower downstream cost.
@@
Under linear probing, shown in Figure 4(a), the most pronounced classwise advantage appears on RGZ DR1: the selected model strengthens all six diagonal entries, reaching 92%, 81%, 78%, 85%, 72%, and 87% from 1C1P through 3C3P. The gain is therefore distributed across the full label space rather than concentrated in a single row. MiraBest is also favorable in this comparison, with the selected model reaching 68% on FR I and 72% on FR II, ahead of both ViT-MAE and DINOv2 (Registers). LoTSS DR2 remains mixed: the selected model leads on FR I, FR II, and Spiral, but the Hybrid and Relaxed-double rows remain strongest for ViT-MAE.
Under full fine-tuning, shown in Figure 4(b), the classwise margins are smaller. On MiraBest, all three models are tied at 70% on FR I, while the selected model matches the strongest FR II row at 75%. LoTSS DR2 remains the most evident limitation: the selected model is strongest only on FR II (88%). RGZ DR1 remains partly favorable, but the gain is more localized than under probing: the selected model matches the best 1C1P row at 93% and leads on 1C2P and 2C2P (90% and 94%), whereas DINOv2 (Registers) remains stronger on 1C3P, 2C3P, and 3C3P. The fine-tuning deltas are therefore associated with selective classwise changes rather than with a uniform shift of the full matrix.
@@ 10.4 Overall Interpretation and Limitations
The branch comparison indicates that additional stages are not uniformly beneficial across settings. Reconstruction-only continued pretraining is the weakest branch and often underperforms the ViT-MAE starting point. Masked reconstruction alone is therefore not sufficient to improve transfer in this setting. Introducing contrastive learning alters the comparison substantially: contrastive-only training already produces large gains over phase 1 alone, especially under linear probing. Across the branch comparison, the two-stage branch is the most consistently top-performing STRADAViT variant, even though a small number of settings—most notably fine-tuned RGZ DR1—are matched or marginally exceeded by contrastive-only training.
@@
Relative to ViT-MAE, the strongest two-stage models improve on all linear-probe settings and on fine-tuned MiraBest and RGZ DR1. Most of this benefit comes from the contrastive objective, while the reconstruction stage contributes a smaller, dataset-dependent benefit and functions primarily as a domain-adaptive warm start for the subsequent contrastive stage. Two factors appear most influential: contrastive losses and ROI-aware view generation. The former better fits a setting in which many negatives are easy background cases while a smaller set are genuinely confusable; the latter makes both contrastive learning and masked reconstruction workable on sparse radio astronomy cutouts. This difference between frozen-feature gains and end-to-end gains is explicit in Table 15 and Figure 4: linear-probe improvements are larger and more consistent across datasets, whereas fine-tuning gains are smaller, more localized to specific classes, and change sign on LoTSS DR2. The classwise comparisons therefore indicate broader representation-level gains under probing than under end-to-end adaptation.
@@
Across the contrastive objectives, InfoNCE provides a baseline but rarely defines the strongest probe results. Soft-HCL is the most consistent choice on LoTSS DR2 and in many register-based settings, whereas HCL tends to give the best MiraBest results and remains strongest on RGZ DR1 fine-tuning. Register-token effects are also more favorable inside the STRADAViT pipeline than in the off-the-shelf DINO baselines, especially on LoTSS DR2 and RGZ DR1 under probing, although the gain is not uniform across every branch and dataset. The DINOv2 initialization ablation is consistent with this picture: it confirms portability of the contrastive adaptation recipe to a stronger starting point, but under the same recipe the adapted register-based DINOv2 initialization becomes preferable to the non-register initialization, despite the opposite ordering in the off-the-shelf linear-probe baselines. The resulting gains remain small and dataset-dependent, with the same LoTSS DR2 limitation that appears across the adapted models more generally.
@@
The view generator is strongly object-centric. This property reduces empty crops and supports training on compact or moderately extended sources, but it also biases the pipeline toward RGZ-style views. LoTSS DR2 and MiraBest contain larger-scale extended structure that benefits from larger-view sampling designed for extended source morphology in radio astronomy. This bias is consistent with the weaker and less uniform LoTSS DR2 outcomes, where extended structure is less well matched by the present view generator.
@@ 11 Conclusions and Future Work
This paper studies how to adapt self-supervised ViT continued pretraining to sparse, heterogeneous radio astronomy imaging. Masked reconstruction alone is insufficient, contrastive learning is the main source of transfer gain, and the two-stage reconstruction-to-contrastive recipe is the most consistent STRADAViT configuration overall. Evidence is strongest under frozen-backbone evaluation, while the fine-tuning results remain favorable in most settings. The targeted DINOv2 ablation shows that the adaptation recipe is not specific to the ViT-MAE starting point, but the ViT-MAE-based two-stage checkpoint remains the selected STRADAViT release because it combines competitive transfer with substantially lower token count and downstream cost than the adapted DINOv2 alternatives. STRADAViT therefore serves as a domain-adapted starting point for radio astronomy vision backbones rather than as a uniformly dominant model on every dataset or class. We release this ViT-MAE-based checkpoint on Hugging Face [^1] https://huggingface.co/ISSA-ML/stradavit-base. Code, configuration files, and dataset-construction scripts are available from the corresponding author upon reasonable request; where direct redistribution of underlying data products is not possible, the reconstruction pipeline and source references can be provided.
Future work should test stronger teacher–student pretraining schemes, especially DINO-style variants built on anchored radio astronomy-aware view generation, together with larger-view regimes that preserve extended emission more consistently in pretraining. Further iterations of the on-the-fly data-curation and view-generation pipeline are also warranted through augmentation and sampling schemes that better match downstream heterogeneity rather than remaining predominantly RGZ-like. The strong linear-probe gains motivate evaluation in fixed-backbone settings, including detection and segmentation pipelines, while the stable downstream gains under full fine-tuning justify evaluation of task-specific models built directly on top of STRADAViT. Additional priorities are broader cross-survey robustness tests, explicit out-of-domain transfer, and repetition of the same controlled pipeline from stronger off-the-shelf initializations.
@@ Acknowledgements
The authors acknowledge Xjenza Malta for funding this study under the STRADA project through the Research Excellence Programme of 2024 [REP-2024-19]. The authors thank Alessio Magro and Kristian Grixti, Institute of Space Sciences and Astronomy, University of Malta, for technical support and for maintaining the GPU computing infrastructure used for the training and evaluation runs in this study.